multiply and write each answer in simplest form
4 1/5 x 10/21 x 9/20

Answer by stanbon(48546)

4 1/5 x 10/21 x 9/20
= 21/5 * 10/21 * 9/20
21's cancel
5 reduces with 10
= 1/1 * 2/1 * 9/20
2 reduces with 20
= 1 * 1* 9/10
= 9/10
============
Cheers,
Stan H.